Session: Nov 1998 (session 4)
Wavelength: 5 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (C. Bortolotti at Sat Nov 14 06:32:20 MET 1998) Success
No known problems.

Onsala (A. Polatidis at Sun Nov 15 12:08:31 MET 1998) Success
Intermitent interference, which lasts for a fraction of a second at a time. It is probably a radar. There was some loss of data (1-2 min at a time) in the scans to and after 3C345, due to long slew time, which the schedule did not allow for. No other problems.

Torun (K.Borkowski at Mon Nov 16 14:05:53 MET 1998) Minor failures
Wrong frequency set on BBC05 until 8:25 UT. 9:14 to 9:24 off source due to long azimuth slew. Later, many times late on source (too short slewing times allowed in the sched).

Effelsberg (dgraham at Mon Nov 16 16:44:57 MET 1998) Failure
Used schedule version 5 instead of version 6 (!) until 1225, so frequencies wrong for 1st 8 hours. Antenna stopped 1658-1807.

Jodrell (Alastair Gunn at Wed Nov 18 11:25:34 MET 1998) Minor failures
Cable wrap at approx 01:04 UT, telescope on source again at approx 01:35 UT. No other known problems.

Calibration (Jean\-Francois Desmurs at Tue Nov 24 19:06:45 MET 1998) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the nov98 directory under the name: gb026.antab.Read the comments in the file to get further information.